OpenStudy (anonymous):

Find the quotient of 3x^3 + 9x^2 -6x / -3x

OpenStudy (anonymous):

Is the quotient : \[\frac{3x^3+9x^2-6x}{-3x}?\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

OK ! Use the rule : \[\frac{a+b+c}{d}=\frac ad+\frac bd+\frac cd\]

OpenStudy (anonymous):

Okay so 3/-3 9/-3 6/-3 ? I honestly have no idea...

OpenStudy (anonymous):

Listen to me ! It is very simple : \[\frac{3x^3+9x^2−6x}{−3x}=\frac{3x^3}{-3x}+\frac{9x^2}{-3x}+\frac{-6x}{-3x} \\~~~~~~~~~~~~~~~~~~~~~~~~=-x^2-3x+2 \]